Grade 12th passPhysical Chemistry

.66kg ball is moving with speed of 100 mtr /sec. The associated wavelength will be.?

Devki Nandan
8 Years agoGrade 12th pass

1 Answer

sonika p
8 Years ago
wavelength is given by de Broglie’s formula, lambda = h/mv
m = 0.66kg, v =100m/s
lambda = (6.636 *10^-34) / (0.66*100) = 1.0039 * 10^ -35 m